OpenStudy (anonymous):

A man earns $112 interest in one year on an investment of $1,400. What was his rate of interest? 8% 10% 12%

OpenStudy (wolf1728):

Here's the formula you need: log(1 + rate) = {log(total) -log(Principal)} ÷ Years log(1 + rate) = log(1,512.00) -log(1,400.00) / 1 log(1 + rate) = (3.1795517912 -3.1492191127) / 1 log(1 + rate) = 0.0303326785 / 1 log(1 + rate) = 0.0303326785 If log (1 + rate) = 0.0303326785 then 10^0.0303326785 = 1 + rate 1+ rate = 1.0723404255 Therefore rate = .0723404255 or .0723404255*100 7.23404255%

jhonyy9 (jhonyy9):

1400 -------------100% 112 -------------- x% ---------------------- x = 112*100/1400 = 112/14 = 8% 8% was his rate of interest

OpenStudy (wolf1728):

Yes it seems I made a mistake in calculating the log of 1,400: log(1 + rate) = (3.1795517912 -3.146128) / 1 log(1 + rate) = 0.0334237555 / 1 log(1 + rate) = 0.0334237555 If log (1 + rate) = 0.0334237555 then 10^0.0334237555 = 1 + rate 1+ rate = 1.08 Therefore rate = .08 or 8%